# TLUE: A Tibetan Language Understanding Evaluation Benchmark

### v2

@@ Abstract
Large language models (LLMs) have made tremendous progress in recent years, but low-resource languages, such as Tibetan, remain significantly underrepresented in their evaluation. Despite Tibetan being spoken by over seven million people, it has largely been neglected in the development and assessment of LLMs. To address this gap, we present TLUE (A Tibetan Language Understanding Evaluation Benchmark), the first large-scale benchmark for assessing LLMs’ capabilities in Tibetan. TLUE comprises two major components: (1) a comprehensive multi-task understanding benchmark spanning 5 domains and 67 subdomains, and (2) a safety benchmark covering 7 subdomains. We evaluate a diverse set of state-of-the-art LLMs. Experimental results demonstrate that most LLMs perform below the random baseline, highlighting the considerable challenges LLMs face in processing Tibetan, a low-resource language. TLUE provides an essential foundation for driving future research and progress in Tibetan language understanding and underscores the need for greater inclusivity in LLM development.
@@ 1 Introduction
Large language models (LLMs) have made significant strides in natural language understanding, excelling in high-resource languages like English and Chinese through extensive datasets, advanced architectures, and benchmarks like GLUE and SuperGLUE Wang et al. (2018, 2019). However, they largely overlook low-resource languages, which limits the inclusivity and fairness of AI systems and restricts their applicability to underserved language communities Bender et al. (2021); Huang et al. (2025). Tibetan, spoken by over seven million people, is a key example of such a language, with unique grammatical features and limited annotated data of Statistics of China (2021). Despite its cultural importance, Tibetan is underrepresented in current LLM research and benchmarks An (2023).
Most language understanding benchmarks, including GLUE, SuperGLUE, and BIG-bench, focus on high-resource languages Wang et al. (2018, 2019); Srivastava et al. (2022), leaving low-resource languages like Tibetan without tailored evaluation frameworks. Existing methods do not adequately capture Tibetan’s linguistic intricacies Liu et al. (2022); Lv et al. (2025). This paper introduces TLUE: A Tibetan Language Understanding Evaluation Benchmark, the first large-scale benchmark for Tibetan, designed to address the unique challenges of low-resource language evaluation.
Our key contributions are as follows:
  • Developed and released TLUE, the first large-scale Tibetan language evaluation benchmark, covering 67 knowledge subdomains (Ti-MMLU) and 7 safety categories (Ti-SafetyBench).
  • Benchmarked leading LLMs (GPT-4O, Claude-Sonnet, Gemini 1.5, LLaMA 3.1, Qwen2.5, Deepseek) to assess their Tibetan language understanding and safety performance.
  • Conducted qualitative analysis identifying significant limitations in current models, with most performing below the random baseline, emphasizing the need for advancements in low-resource language modeling.
@@ 2 Related Work
Language Understanding Benchmarks:
LLMs have been extensively evaluated on multilingual and domain-specific benchmarks, such as XTREME Hu et al. (2020), XGLUE Liang et al. (2020), and MASSIVE FitzGerald et al. (2022). However, these benchmarks predominantly focus on high-resource languages, with little attention given to low-resource languages like Tibetan.
Low-Resource Language Evaluation:
Several efforts have been made to extend LLM evaluation to low-resource languages. IndicGLUE Kakwani et al. (2020), and INDICGENBENCH Singh et al. (2024) focus on Indic languages, while projects such as AmericasNLP Mager et al. (2021) and Masakhane $\forall$ et al. (2020) target indigenous and African languages. Despite these advances, Tibetan remains significantly underrepresented in LLM evaluation. To address this issue, we propose Ti-MMLU, a large-scale Tibetan understanding benchmark inspired by the task design of CMMLU Li et al. (2023), enabling systematic evaluation of model performance in a challenging low-resource setting.
Safety and Ethical Considerations in LLMs:
Recent work has examined the safety and ethical risks of LLMs, particularly in multilingual and low-resource contexts. Studies such as SafetyBench Zhang et al. (2023) and HolisticEval Liang et al. (2023) assess biases, fairness, and robustness in AI systems. However, these benchmarks largely exclude Tibetan and other low-resource languages, leaving significant gaps in understanding how safety concerns manifest in such linguistic environments. Ti-SafetyBench introduces a dedicated safety evaluation suite for Tibetan, providing insights into model robustness and potential vulnerabilities in a low-resource language context.
In summary, the combination of Ti-MMLU and Ti-SafetyBench forms TLUE, the first large-scale Tibetan language understanding evaluation benchmark in the world. By offering a comprehensive, multi-domain framework for both language understanding and safety evaluation, TLUE fills a critical gap in the current landscape of LLM benchmarks, enabling systematic assessment of Tibetan language models in a challenging low-resource setting.
@@ 3 TLUE
In designing TLUE, we adhere to several key principles to ensure a rigorous and meaningful evaluation of Tibetan LLMs. First, we align our benchmark with existing multilingual evaluation frameworks to facilitate direct comparison with established benchmarks. This allows researchers to assess the performance of Tibetan LLMs relative to models trained in other languages, providing a broader perspective on cross-lingual capabilities Park et al. (2024). Additionally, we ensure that the evaluation data is carefully curated and free from data leakage, maintaining the integrity of benchmark results and preventing unintended memorization effects Doe and Smith (2024).
Figure 2: The curation process of TLUE
Curation process.
In the curation process outlined in Figure 2, the TLUE benchmark follows a structured multi-stage process to ensure the quality and cultural relevance of Tibetan evaluation data. Source datasets, such as CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), are first translated into Tibetan using Claude 3.5 Sonnet Anthropic (2024), which demonstrates strong Tibetan text generation capabilities. The translations undergo rule-based filtering to correct syntactic inconsistencies and structural errors.
Following this, expert human annotators, including two Tibetan language specialists and a team of 15 annotators, refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. Annotators are compensated at an hourly rate of 28 USD, ensuring high-quality review and incentivizing skilled professionals.
@@
Dataset Sizes.
@@
TLUE consists of 22,963 evaluation questions, divided into Ti-MMLU for knowledge assessment and Ti-SafetyBench for safety evaluation. Ti-MMLU includes 11,528 multiple-choice questions across 67 subjects, covering social sciences, STEM, China-specific topics, humanities, and other domains, enabling a comprehensive evaluation of Tibetan language understanding. Ti-SafetyBench contains 11,435 questions across 7 safety categories, including ethics, bias, health, and privacy, designed to assess LLMs’ ability to handle sensitive content. Dataset sizes and licenses are detailed in Appendix Table 3, adhering to the original dataset licenses where applicable.While Ti-MMLU focuses on broad, multi-domain knowledge evaluation, Ti-SafetyBench targets high-precision safety assessments, together forming a balanced and comprehensive Tibetan language benchmark. The dataset sizes are summarized in Figure 1, which illustrates the comprehensive breakdown of Ti-MMLU and Ti-SafetyBench, highlighting the wide range of subjects and safety categories covered.
Why Extend Existing Benchmarks?
@@
Developing a benchmark for Tibetan, a low-resource language, presents challenges such as limited annotated corpora and linguistic expertise. TLUE extends existing benchmarks via translation, enhanced with human annotation and quality control. Key reasons include:
@@
  • Enabling Cross-Lingual Comparability: A translation-based benchmark provides parallel data to compare Tibetan model performance across languages, isolating linguistic effects from task knowledge.
@@
  • Addressing Data Scarcity and Leveraging Established Designs: Due to limited Tibetan corpora, TLUE adapts reliable, pre-validated benchmarks like CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), ensuring task diversity and alignment with Tibetan’s linguistic and cultural traits.
@@
  • Ensuring Quality through Controlled Adaptation: TLUE uses a multi-stage adaptation process (translation, filtering, alignment, annotation) to balance efficiency with linguistic precision, maintaining high benchmark standards.
This structured approach provides a high-quality, scalable framework for assessing LLM performance in Tibetan, especially in low-resource settings.
4 Experiments and Analysis
@@
We use TLUE to evaluate the Tibetan language understanding and safety capabilities of various LLMs. Our experiments cover a diverse range of models, including both open-source and proprietary models.
@@
  • Open-source models. We evaluate several open-source LLMs, including LLaMA-3.1-405B-Instruct Dubey et al. (2024), LLaMA-3.1-8B-Instruct Dubey et al. (2024), Qwen2.5-72B-Instruct Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024), and DeepSeek-R1 DeepSeek-AI et al. (2025). These models represent a variety of architectures and parameter scales, allowing for a broad comparison of open-source advancements in Tibetan language understanding.
@@
  • Proprietary models. We also evaluate several proprietary models, including GPT-4OOpenAI (2024), GPT-3.5-Turbo OpenAI (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Flash-Latest DeepMind (2024), and O1-Mini OpenAI (2024). These models serve as strong baselines for commercial LLM performance in Tibetan tasks.
We investigate several key aspects of LLM performance in Tibetan. We first evaluate all models on Ti-MMLU and Ti-SafetyBench, measuring their Tibetan multi-task understanding and safety capabilities. We then assess the impact of language resource availability by comparing CMMLU (Chinese) and Ti-MMLU(Tibetan) and extend this analysis to safety evaluation by comparing SafetyBench (Chinese) with Ti-SafetyBench. Furthermore, we explore the differences between reasoning-optimized models and chat models, examining whether reasoning enhancements improve low-resource language performance. Finally, we study the effect of model size on Tibetan language understanding, comparing different parameter variants of LLaMA-3.1 and Qwen2.5.
Evaluation Methodology.
@@
We employ zero-shot evaluation to simulate real-world conditions where models must perform tasks without task-specific examples, providing a more accurate assessment of performance in low-resource languages like Tibetan. Appendix Figure 6 illustrates the TLUE prompt format and example queries used in evaluation. To ensure consistency, we use the default temperature settings for each model during inference, avoiding bias from manual adjustments.
Evaluation Metrics.
@@
We evaluate model performance on TLUE using two metrics: response rate and accuracy. Response rate measures the proportion of valid responses, while accuracy tracks the proportion of correct answers. For Ti-MMLU, we also compute conditional accuracy, which reflects the model’s performance when giving valid responses (see Appendix). We use two methods for metric calculation:
  • Direct Answer Calculation (DA): Measures if the model selects a valid and correct answer. See Appendix Algorithm 1 for details.
@@
  • Concern All Answer Calculation (CAA): Evaluates the model’s maximum performance by considering all valid options and selecting the correct answer based on remaining choices. See Appendix Algorithm 2 for details.
These methods provide a thorough assessment of the model’s Tibetan language performance, with DA focusing on correct responses and CAA on the model’s potential across all choices.
4.1 Overall Performance on TLUE
Model Eval Method Avg. STEM Humanities Social Sciences Other China specific
Claude-3.5-Sonnet DA 33.95 28.04 36.32 39.31 34.01 32.09
Gemini-1.5-Flash 30.14 25.75 29.72 35.9 29.61 29.7
@@ 4.1 Overall Performance on TLUE
GPT-3.5-Turbo 2.11 2.40 2.18 2.20 1.88 1.87
Deepseek-v3 29.51 23.57 31.97 33.65 29.92 28.44
Llama3.1-405B-Instruct 25.08 23.88 24.25 25.58 27.62 24.07
Llama3.1-70B-Instruct 23.73 23.16 23.2 26.2 24.65 21.45
Qwen2.5-32B-Instruct 13.94 12.63 14.98 14.92 13.71 13.44
Qwen2.5-72B-Instruct 7.27 6.07 7.98 7.52 7.74 7.02
Qwen2.5-7B-Instruct 1.8 2.94 1.87 1.63 0.9 1.68
Llama3.1-8B-Instruct 5.47 5.48 5.56 5.99 5.46 4.86
Deepseek-r1 15.74 13.87 14.24 18.06 13.94 18.58
O1-mini 6.14 6.15 6.33 7.17 6.06 4.98
Claude-3.5-Sonnet CAA 35.63 30.88 37.47 40.58 35.26 33.96
Gemini-1.5-Flash 31.01 26.68 30.58 36.69 30.30 30.81
@@
GPT-3.5-Turbo 3.42 3.82 3.35 3.68 3.09 3.16
Deepseek-v3 32.16 27.03 34.58 36.26 32.00 30.94
Llama3.1-405B-Instruct 25.28 24.10 24.50 25.87 27.73 24.22
Llama3.1-70B-Instruct 23.79 23.22 23.24 26.31 24.65 21.52
Qwen2.5-32B-Instruct 18.56 16.66 20.3 19.72 17.47 18.67
Qwen2.5-72B-Instruct 16.50 15.73 17.88 17.00 15.84 16.04
Qwen2.5-7B-Instruct 14.59 13.92 13.66 16.34 14.57 14.46
Llama3.1-8B-Instruct 7.44 7.95 7.54 7.38 7.41 6.92
Deepseek-r1 27.45 21.01 25.99 32.72 26.44 31.08
O1-mini 9.67 9.69 9.80 10.14 9.68 9.02
@@
Table 1: Accuracy results of different LLMs on the Ti-MMLU dataset using two evaluation methods: Concern All Answer (CAA) and Direct Answer (DA).
Tables 1 and 2 show results on Ti-MMLU and Ti-SafetyBench, evaluating Tibetan language understanding and safety.The detial data is released in Appendix Table 14, 15, 16, 17
Performance on Ti-MMLU.

### v3

@@ Abstract
Large language models have made tremendous progress in recent years, but low-resource languages, like Tibetan, remain significantly underrepresented in their evaluation. Despite Tibetan being spoken by over seven million people, it has largely been neglected in the development and assessment of LLMs. To address this gap, we present a Tibetan Language Understanding Evaluation Benchmark, TLUE, which is also the first large-scale benchmark for measuring the proficiency of large language models in the Tibetan language. TLUE comprises two major components: a comprehensive multi-task understanding benchmark spanning 5 domains and 67 subdomains, and a safety benchmark encompassing 7 subdomains. Finally, we evaluate a diverse set of state-of-the-art LLMs. Experimental results demonstrate that most large language models perform below the random baseline, especially highlighting the considerable challenges they face in Tibetan language processing. TLUE provides a crucial foundation for advancing future research in Tibetan language understanding and highlights the importance of promoting greater inclusivity in the development of large language models.
@@ 1 Introduction
Large language models (LLMs) have made significant strides in natural language understanding, excelling in high-resource languages like English and Chinese through extensive datasets, advanced architectures, and benchmarks like GLUE Wang et al. (2018) and SuperGLUE Wang et al. (2019). However, many mainstream LLMs, like GPT OpenAI (2024); Achiam et al. (2023), Claude Anthropic (2024), Gemini DeepMind (2024), LlaMA Dubey et al. (2024), Qwen Team (2024) and DeepSeek DeepSeek-AI et al. (2025, 2024) largely overlook low-resource languages, which limits the inclusivity and fairness of AI systems and restricts their applicability to underserved language communities Bender et al. (2021); Huang et al. (2025). Tibetan is one of the languages included, spoken by over seven million people. It has unique grammatical features and limited annotated data of Statistics of China (2021). Despite its cultural importance, Tibetan is underrepresented in current LLM research and benchmarks An (2023); Liu et al. (2025).
Current language understanding benchmarks, including GLUE Wang et al. (2018), SuperGLUE Wang et al. (2019), and BIG-bench Srivastava et al. (2022), focus on high-resource languages, leaving low-resource languages like Tibetan without tailored evaluation frameworks, including benchmarks from China Li et al. (2023); Zhang et al. (2023). Even existing methods do not adequately capture Tibetan’s linguistic intricacies Liu et al. (2022); Lv et al. (2025). In this paper, we introduce the TLUE, a Tibetan Language Understanding Evaluation benchmark, the first large-scale benchmark for Tibetan, designed to address the unique challenges of low-resource language evaluation. Furthermore, we selected several state-of-the-art large language models for training and evaluation, and the experimental results demonstrate that TLUE effectively exposes the limitations of current models in handling Tibetan language tasks. Most LLMs show significantly lower performance on TLUE compared to benchmarks in high-resource languages, particularly struggling with tasks requiring deep linguistic understanding and domain-specific knowledge. This highlights the urgent need for more inclusive pretraining data and architecture adaptations tailored to low-resource languages like Tibetan. Our findings validate TLUE as a robust tool for evaluating and guiding the development of more equitable language models.
The main contributions of our work are summarized as follows:
  • We developed and publicly released TLUE, the first large-scale Tibetan Language Understanding Evaluation benchmark. It is designed to fill the gap in existing resources by providing a comprehensive suite of tasks that span 67 knowledge-based subdomains and 7 safety-critical categories, addressing both general understanding and ethical alignment in Tibetan.
  • We evaluated several state-of-the-art large language models, including GPT-4O Achiam et al. (2023) and GPT-O1-mini OpenAI (2024), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5 DeepMind (2024), LlaMA-3.1 Dubey et al. (2024), Qwen-2.5 Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024) and DeepSeek-R1 DeepSeek-AI et al. (2025), to systematically assess their capabilities in both general Tibetan language understanding and safety-oriented tasks. These LLMs represent a diverse set of architectures and training strategies, enabling a broad comparison of their performance under low-resource conditions.
  • Our qualitative and quantitative analyses reveal substantial limitations in current large language models when applied to Tibetan. Most models performed significantly below expected baselines, even falling below random choice levels on certain tasks, particularly in domains requiring complex reasoning or cultural sensitivity. These findings underscore the urgent need for research into more inclusive, linguistically-aware model development tailored to low-resource languages.
@@ 2 Related Work
2.1 Language Understanding Benchmark
LLMs have been extensively evaluated on multilingual and domain-specific benchmarks, such as XTREME Hu et al. (2020), XGLUE Liang et al. (2020), and MASSIVE FitzGerald et al. (2022). However, these benchmarks predominantly focus on high-resource languages, with little attention given to low-resource languages like Tibetan. CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023) introduced large-scale Chinese language evaluations, but no equivalent benchmark existed for Tibetan.
2.2 Low Resource Language Evaluation
Several efforts have been made to extend LLM evaluation to low-resource languages. IndicGLUE Kakwani et al. (2020), and INDICGENBENCH Singh et al. (2024) focus on Indic languages, while projects such as AmericasNLP Mager et al. (2021) and Masakhane $\forall$ et al. (2020) target indigenous and African languages. Despite these advances, Tibetan remains significantly underrepresented in LLM evaluation.
2.3 Safety and Ethical Limitations in LLM
Recent work has also examined the safety and ethical risks of LLMs, particularly in multilingual and low-resource contexts. Studies such as SafetyBench Zhang et al. (2023) and HolisticEval Liang et al. (2023) assess biases, fairness, and robustness in AI systems. However, these benchmarks largely exclude Tibetan and other low-resource languages, leaving significant gaps in understanding how safety concerns manifest in such linguistic environments.
2.4 Motivation
Based on the current progress of LLM in minority languages, we propose the TLUE, filling the gap by providing a comprehensive Tibetan evaluation dataset across multiple domains and safety-related tasks. We referred to the two benchmarks, CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), and asked ancient Tibetan experts to manually translate and verify them. In summary, the combination of Ti-MMLU and Ti-SafetyBench forms TLUE. TLUE can not only enable systematic evaluation of model performance in a challenging low-resource setting, but also introduces a dedicated safety evaluation suite for Tibetan, providing insights into model robustness and potential vulnerabilities in a low-resource language context.
By offering a comprehensive, multi-domain framework for both language understanding and safety evaluation, TLUE fills a critical gap in the current landscape of LLM benchmarks, enabling systematic assessment of Tibetan language models in a challenging low-resource setting.
Figure 1: The Curation Process of TLUE
@@ 3 TLUE
In designing TLUE, we adhere to several key principles to ensure a rigorous and meaningful evaluation of Tibetan LLMs.
First, we align our benchmark with existing multilingual evaluation frameworks to facilitate direct comparison with established benchmarks. This allows researchers to assess the performance of Tibetan LLMs relative to models trained in other languages, providing a broader perspective on cross-lingual capabilities Park et al. (2024).
Additionally, we ensure that the evaluation data is carefully curated and free from data leakage, maintaining the integrity of benchmark results and preventing unintended memorization effects Doe and Smith (2024).
3.1 Curation Process
As shown in Figure 1, the TLUE benchmark follows a structured multi-stage process to ensure the quality and cultural relevance of Tibetan evaluation data in the curation process. Source datasets, such as CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), are first translated into Tibetan using Claude-3.5-Sonnet Anthropic (2024) , which demonstrates strong Tibetan text generation capabilities. The translations undergo rule-based filtering to correct syntactic inconsistencies and structural errors.
Following this, expert human annotators, including 2 Tibetan language specialists and their team of 5 annotators, refine the dataset to preserve linguistic accuracy and classical Tibetan grammar. Annotators are compensated at an hourly rate of 28 USD, ensuring high-quality review and incentivizing skilled professionals.
@@ 3.1 Curation Process
3.2 Size
TLUE consists of 22,963 evaluation questions, divided into Ti-MMLU for knowledge assessment and Ti-SafetyBench for safety evaluation.
@@ 3.2 Size
Ti-MMLU includes 11,528 multiple-choice questions across 67 subjects, covering social sciences, STEM, China-specific topics, humanities, and other domains, enabling a comprehensive evaluation of Tibetan language understanding. Ti-SafetyBench contains 11,435 questions across 7 safety categories, including ethics, bias, health, and privacy, designed to assess LLMs’ ability to handle sensitive content.
While Ti-MMLU focuses on broad, multi-domain knowledge evaluation, Ti-SafetyBench targets high-precision safety assessments, together forming a balanced and comprehensive Tibetan language benchmark. The dataset sizes are summarized in Appendix A Figure 5, which illustrates the comprehensive breakdown of Ti-MMLU and Ti-SafetyBench, highlighting the wide range of subjects and safety categories covered. For more details of such two sub-benchmarks, they are shown in Appendix A.
3.3 Why Extend Existing Benchmarks?
@@
Developing a benchmark for Tibetan presents challenges such as limited annotated corpora and linguistic expertise. TLUE extends existing benchmarks via translation, enhanced with human annotation and quality control. Key reasons include:
3.3.1 Enabling Cross-Lingual Comparability
@@
A translation-based benchmark provides parallel data to compare Tibetan model performance across languages, isolating linguistic effects from task knowledge.
3.3.2 Addressing Data Scarcity and Leveraging Established Designs
@@
Due to limited Tibetan corpora, TLUE adapts reliable, pre-validated benchmarks like CMMLU Li et al. (2023) and SafetyBench Zhang et al. (2023), ensuring task diversity and alignment with Tibetan’s linguistic and cultural traits.
3.3.3 Ensuring Quality through Controlled Adaptation
@@
TLUE uses a multi-stage adaptation process (translation, filtering, alignment, annotation) to balance efficiency with linguistic precision, maintaining high benchmark standards. This structured approach provides a high-quality, scalable framework for assessing LLM performance in Tibetan, especially in low-resource settings.
4 Implementation
@@
Our experiments cover a diverse range of LLMs, including both open-source and proprietary LLMs.
@@
For Open-source LLM, we evaluate several open-source LLMs, including LlaMA-3.1-405B Dubey et al. (2024), LlaMA-3.1-8B Dubey et al. (2024), Qwen-2.5-72B Team (2024), DeepSeek-V3 DeepSeek-AI et al. (2024), and DeepSeek-R1 DeepSeek-AI et al. (2025). These LLMs represent a variety of architectures and parameter scales, allowing for a broad comparison of open-source advancements in Tibetan language understanding.
@@
For Proprietary LLM, we also evaluate several proprietary LLMs, including GPT-4O OpenAI (2024), GPT-3.5-Turbo Achiam et al. (2023), Claude-3.5-Sonnet Anthropic (2024), Gemini-1.5-Flash-Latest DeepMind (2024), and O1-Mini OpenAI (2024). These models serve as strong baselines for commercial LLM performance in Tibetan tasks.
4.1 Experimental Setup
We investigate several key aspects of LLM performance in Tibetan:
$\bullet$ We evaluate all LLMs on Ti-MMLU and Ti-SafetyBench, measuring their Tibetan multi-task understanding and safety capabilities.
$\bullet$ We assess the impact of language resource availability by comparing CMMLU Li et al. (2023) and Ti-MMLU and extend this analysis to safety evaluation by comparing SafetyBench Zhang et al. (2023) with Ti-SafetyBench.
$\bullet$ We explore the differences between reasoning-optimized LLMs and chat LLMs, examining whether reasoning enhancements improve low-resource language performance.
$\bullet$ We study the effect of LLM size on Tibetan language understanding, comparing different parameter variants of LlaMA-3.1 Dubey et al. (2024) and Qwen-2.5 Team (2024).
4.2 Evaluation
4.2.1 Evaluation Methodology
@@
We employ zero-shot evaluation to simulate real-world conditions where models must perform tasks without task-specific examples, providing a more accurate assessment of performance in Tibetan. Appendix Figure 7 illustrates the TLUE prompt format and example queries used in evaluation. To ensure consistency, we use the default temperature settings for each model during inference, avoiding bias from manual adjustments.
4.2.2 Evaluation Metrics
@@
We evaluate LLM performance on TLUE using two metrics: response rate and accuracy. Response Rate measures the proportion of valid responses, while Accuracy tracks the proportion of correct answers. We also compute Conditional Accuracy, which reflects the LLM’s performance when giving valid responses.
For both Ti-MMLU and Ti-SafetyBench, We use two methods for metric calculation:
  • Direct Answer Calculation measures if the LLM selects a valid and correct answer.
@@
  • Concern All Answer Calculation evaluates the LLM’s maximum performance by considering all valid options and selecting correct answers based on remaining choices.
These methods provide a thorough assessment of LLMs’ Tibetan language performance. For more details of Ti-MMLU and Ti-SafetyBench, please refer to Appendix B Algorithm 1 and Appendix B Algorithm 2.
5 Performance on TLUE
This section shows the performance of LLM on the TLUE benchmark. We use abbreviated forms for some of the special names or evaluation matrix names. Tables 1(a) and 1(b) show results on Ti-MMLU and Ti-SafetyBench, evaluating Tibetan language understanding and safety.
Method LLM Version Avg. STEM Human Social Other China
DA Claude 3.5-Sonnet 33.95 28.04 36.32 39.31 34.01 32.09
Gemini 1.5-Flash 30.14 25.75 29.72 35.9 29.61 29.7
DeepSeek R1 15.74 13.87 14.24 18.06 13.94 18.58
DeepSeek V3 29.51 23.57 31.97 33.65 29.92 28.44
@@ 5 Performance on TLUE
GPT 3.5-Turbo 2.11 2.40 2.18 2.20 1.88 1.87
GPT O1-mini 6.14 6.15 6.33 7.17 6.06 4.98
Qwen 2.5-32B 13.94 12.63 14.98 14.92 13.71 13.44
Qwen 2.5-72B 7.27 6.07 7.98 7.52 7.74 7.02
Qwen 2.5-7B 1.8 2.94 1.87 1.63 0.9 1.68
LlaMA 3.1-405B 25.08 23.88 24.25 25.58 27.62 24.07
LlaMA 3.1-70B 23.73 23.16 23.2 26.2 24.65 21.45
LlaMA 3.1-8B 5.47 5.48 5.56 5.99 5.46 4.86
CAA Claude 3.5-Sonnet 35.63 30.88 37.47 40.58 35.26 33.96
Gemini 1.5-Flash 31.01 26.68 30.58 36.69 30.30 30.81
DeepSeek R1 27.45 21.01 25.99 32.72 26.44 31.08
DeepSeek V3 32.16 27.03 34.58 36.26 32.00 30.94
@@
GPT 3.5-Turbo 3.42 3.82 3.35 3.68 3.09 3.16
GPT O1-mini 9.67 9.69 9.80 10.14 9.68 9.02
Qwen 2.5-32B 18.56 16.66 20.3 19.72 17.47 18.67
Qwen 2.5-72B 16.50 15.73 17.88 17.00 15.84 16.04
Qwen 2.5-7B 14.59 13.92 13.66 16.34 14.57 14.46
LlaMA 3.1-405B 25.28 24.10 24.50 25.87 27.73 24.22
LlaMA 3.1-70B 23.79 23.22 23.24 26.31 24.65 21.52
LlaMA 3.1-8B 7.44 7.95 7.54 7.38 7.41 6.92
@@
(a) Accuracy Performance on the Ti-MMLU
Method LLM Version Avg. OFF UB PH MH IA EM PP
DA Claude 3.5-Sonnet 50.5 32.3 56.1 63.7 61.0 31.7 52.6 65.6
Gemini 1.5-Flash 43.7 32.7 44.2 59.1 49.4 37.6 43.6 46.3
DeepSeek R1 24.3 20.5 37.2 25.8 20.4 16.9 29.2 16.7
DeepSeek V3 37.4 25.5 38.1 40.5 39.1 44.3 38.3 36.9
GPT 4O 31.1 24.0 28.8 24.6 42.1 39.7 22.4 37.6
GPT 3.5-turbo 9.4 8.6 12.3 9.7 8.7 7.3 8.5 10.6
GPT O1-mini 10.9 11.1 17.0 9.7 9.5 7.0 10.9 9.5
Qwen 2.5-72B 21.2 19.8 44.1 15.0 16.4 11.8 19.9 15.9
Qwen 2.5-32B 21.9 19.9 37.6 17.6 20.3 18.1 17.8 18.6
Qwen 2.5-7B 9.0 10.0 16.8 7.3 5.7 6.6 7.7 7.2
LlaMA 3.1-405B 43.5 36.8 31.4 46.0 52.2 50.4 44.8 46.2
LlaMA 3.1-70B 37.0 32.2 37.3 30.0 40.9 44.0 34.2 39.1
LlaMA 3.1-8B 9.9 9.7 10.9 9.4 10.1 8.8 10.5 9.9
CAA Claude 3.5-Sonnet 58.5 51.1 56.5 66.1 66.9 50.1 57.9 67.6
Gemini 1.5-Flash 49.6 44.2 44.3 60.2 56.0 44.9 51.6 51.0
DeepSeek R1 46.8 42.9 45.7 51.1 50.0 45.7 55.8 33.9
DeepSeek V3 48.3 44.3 44.9 51.1 46.4 55.6 51.8 43.6
GPT 4O 32.9 28.7 30.1 25.5 42.9 40.9 24.8 38.7
GPT 3.5-turbo 11.6 12.5 16.0 11.4 10.3 8.2 10.8 11.8
GPT O1-mini 15.1 16.3 23.1 13.5 13.4 9.9 15.4 11.7
Qwen 2.5-72B 34.1 34.8 51.6 30.9 30.7 25.5 31.3 30.6
Qwen 2.5-32B 30.6 36.0 45.2 28.5 24.4 22.6 28.6 24.9
Qwen 2.5-7B 30.2 35.2 39.3 27.1 25.9 23.2 31.2 25.8
LlaMA 3.1-405B 43.9 37.6 31.7 46.5 52.4 50.8 45.1 46.4
LlaMA 3.1-70B 37.4 32.9 37.3 30.5 41.1 44.5 35.0 39.3
LlaMA 3.1-8B 12.0 12.4 12.6 11.9 12.3 10.1 13.0 11.5
- Random - 36.7 34.5 49.9 27.6 49.5 28.0 26.0 36.4
(b) Accuracy Performance on the Ti-SafetyBench
Table 1: Accuracy Performance of LLMs on the TLUE based on CAA and DA ($\times$100%)
5.1 Performance on Ti-MMLU